# Modeling $(\infty,1)$-categories with Segal spaces

### v1

@@ 5. Properties of the model structure
In this section, we show that the categorical model structure $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ constructed in Theorem 3.1 has desirable properties. Specifically, in Section 5.1, we prove that it is cartesian closed, and in Section 5.2, we establish that it is left proper.
@@ 5.1. Cartesian closedness
The model category $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ from Theorem 3.1 is cartesian closed.
@@
By Proposition 3.8, the $\mathcal{I}$-cofibrations $f\colon X\to Y$ and $f^{\prime}\colon X^{\prime}\to Y^{\prime}$ are monomorphisms and there are sets $R$ and $R^{\prime}$ and weak equivalences $X_{0}\amalg R\xrightarrow{\simeq}Y_{0}$ and $X^{\prime}_{0}\amalg R^{\prime}\xrightarrow{\simeq}Y^{\prime}_{0}$ in $\mathrm{sSet}$ whose restrictions to $X_{0}$ and $X^{\prime}_{0}$ are $f_{0}$ and $f^{\prime}_{0}$, respectively. We show that the pushout-product map $f\,\square\,f^{\prime}$ is an $\mathcal{I}$-cofibration by verifying the conditions from Proposition 3.8.
Since the maps $f$ and $f^{\prime}$ are monomorphisms, so is their pushout-product $f\,\square\,f^{\prime}$. Using that products commute with coproducts and that $\mathrm{sSet}$ is cartesian closed, the induced maps
$\left(X_{0}\times X^{\prime}_{0}\amalg X_{0}\times R^{\prime}\amalg R\times X^{\prime}_{0}\right)\amalg R\times R^{\prime}\xrightarrow{\simeq}\left(X_{0}\times Y^{\prime}_{0}\amalg_{X_{0}\times X^{\prime}_{0}}Y_{0}\times X^{\prime}_{0}\right)\amalg R\times R^{\prime}\to Y_{0}\times Y^{\prime}_{0}$
are such that the first map and the composite are weak equivalences in $\mathrm{sSet}$. Hence, by $2$-out-of-$3$, we get that the second map is a weak equivalence in $\mathrm{sSet}$, whose restriction to $X_{0}\times Y^{\prime}_{0}\amalg_{X_{0}\times X^{\prime}_{0}}Y_{0}\times X^{\prime}_{0}$ is $(f\,\square\,f^{\prime})_{0}$. This shows that $f\,\square\,f^{\prime}$ is an $\mathcal{I}$-cofibration. ∎
We now further show that, if one of the two $\mathcal{I}$-cofibrations is a $\mathcal{J}$-cofibration, then their pushout-product is in fact a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, concluding the proof of Theorem 5.1.
@@
We first deal with the case where $f^{\prime}$ is a $\mathcal{J}_{0}$-cofibration, and hence a trivial cofibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ by Remark 3.10. By Proposition 3.8, the $\mathcal{I}$-cofibration $f\colon X\to Y$ is a monomorphism, and so a cofibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$. Since $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ is cartesian closed, we get that the pushout-product map $f\,\square\,f^{\prime}$ is weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$. By $2$-out-of-$3$, a $\mathcal{J}_{0}$-fibrant replacement of $f\,\square\,f^{\prime}$ is also a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ between Segal spaces, and so a Dwyer–Kan equivalence by Example 2.12. This shows that $f\,\square\,f^{\prime}$ is a weak equivalence.
We now treat the case where $f^{\prime}$ is the inclusion $F[0]\hookrightarrow NI[1]$. Showing that the $\mathcal{I}$-cofibration $X\times NI[1]\amalg_{X}Y\to Y\times NI[1]$ is a weak equivalence is equivalent to showing that, for every $\mathcal{J}$-fibration $W\to Z$ between $\mathcal{J}$-fibrant objects, the induced map
@@
is an $\mathcal{I}$-fibration. We prove this by showing that it is both an isofibration and a Dwyer–Kan equivalence. Since $W\to Z$ is an isofibration between Segal spaces, then Lemma 2.18 shows that the desired map is an isofibration. Moreover, by Lemmas 2.18 and 3.16, the maps $W^{NI[1]}\to W$ and $Z^{NI[1]}\to Z$ are both isofibrations and Dwyer–Kan equivalences; hence they are $\mathcal{I}$-fibrations. Then, in the composite
@@
the second map is an $\mathcal{I}$-fibration as the pullback of $Z^{NI[1]}\to Z$ and the composite is a Dwyer–Kan equivalence. Hence, the first map is a Dwyer–Kan equivalence as well, by $2$-out-of-$3$. ∎
@@ 5.2. Left properness
The model category $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ from Theorem 3.1 is left proper.
@@
Let $i\colon A\to B$ be an $\mathcal{I}$-cofibration and $f\colon A\to C$ be a weak equivalence $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. Consider the pushout $g\colon B\to D$ of $f$ along $i$; we want to prove that $g$ is also a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. First note that, without loss of generality, we can assume that $C$ is a Segal space, as $\mathcal{J}$-cofibrations are closed under pushouts and weak equivalences in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ satisfy $2$-out-of-$3$.
@@
We need to check that $qk$ is a weak equivalence. Note that $k$ is a $\mathcal{J}$-cofibration, and hence a weak equivalence, as it is the pushout of the $\mathcal{J}$-cofibration $j$. To see that $q$ is a weak equivalence, note that the pushout $i^{\prime}$ of $i$ is an $\mathcal{I}$-cofibration, so in particular a monomorphism by Proposition 3.8. Since $f$ and $j$ are both weak equivalences in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ and $C$ is a Segal space, the $2$-out-of-$3$ property implies that the $\mathcal{J}$-fibration $p$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ between Segal spaces, and hence a Dwyer–Kan equivalence. By Theorem 4.2 (iii), the map $p$ is then a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. As the model structure $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ is left proper, the pushout $q$ of $p$ along the monomorphism $i^{\prime}$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. By Lemma 4.3, the map $q$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, so that the composite $qk$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ as well. ∎
@@ 6. Applications
In this last section, we present two notable consequences of the existence of the categorical model structure. In Section 6.1, we show that the inclusion of categories into $(\infty,1)$-categories has a particularly simple description in terms of the model category $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. Specifically, we prove that the discrete nerve functor is right Quillen. Then, in Section 6.2, we show that homotopy limits of $(\infty,1)$-categories can be computed in a manner similar to homotopy limits of ordinary categories. Specifically, we prove that the limit of an injectively fibrant diagram for $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ is already a homotopy limit for the complete Segal space model structure.
@@ 6.1. Nerves of categories
The adjunction from Remark 2.7
[…]
@@
Any map of sets is a fibration in $\mathrm{sSet}$, so that any map between simplicial sets (viewed as simplicial spaces) is a Reedy fibration. Since $N\mathcal{C}$ satisfies the Segal conditions, for every category $\mathcal{C}$, and the nerve preserves isofibrations, it follows that $N$ preserves fibrations. Next, the trivial fibrations $f\colon\mathcal{C}\to\mathcal{D}$ in $\mathrm{Cat}$ are the functors that are fully faithful and surjective on objects, so that the induced map $Nf\colon N\mathcal{C}\to N\mathcal{D}$ between the nerves satisfies the equivalent conditions of Lemma 2.11. By Proposition 3.17, the map $Nf$ is then an $\mathcal{I}$-fibration. This implies that $N$ is a right Quillen functor.
Finally, the nerve $N\mathcal{C}$ of any category $\mathcal{C}$ is cofibrant by Proposition 3.8, so that the derived counit is equivalent to the counit. The latter is an isomorphism as the nerve is fully faithful. ∎
@@ 6.2. Homotopy limits of $(\infty,1)$-categories
In fact, this is a special instance of the following result for more general homotopy limits.
@@
Let $\mathcal{C}$ be a small category and consider a diagram $F\colon\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ that is injectively fibrant as a diagram $\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. Then the limit of $F$ is a homotopy limit of $F$ in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$.
@@
Consider an injective fibrant replacement $\alpha\colon F\Rightarrow\widehat{F}$ in the model structure of diagrams $\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. In particular, for every object $c\in\mathcal{C}$, the induced map $\alpha_{c}\colon F(c)\to\widehat{F}(c)$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ between Segal spaces, and so a Dwyer–Kan equivalence by Theorem 4.2 (iii). As $\mathrm{id}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ is a right Quillen functor, the diagram $\widehat{F}\colon\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ is also injectively fibrant as a diagram $\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. Hence, we have a levelwise weak equivalence $\alpha\colon F\Rightarrow\hat{F}$ between injectively fibrant diagrams in the model structure of diagrams $\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. As a consequence, the induced map $\mathrm{lim}_{\mathcal{C}}F\to\mathrm{lim}_{\mathcal{C}}\widehat{F}$ is a Dwyer-Kan equivalence between Segal spaces, and so a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ by Theorem 4.2 (iii). This concludes the proof. ∎
@@
Since $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ is cartesian closed by Theorem 5.1, every Segal space $X$ has a canonical Reedy fibrant simplicial resolution $\Delta^{\mathrm{op}}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ sending $[m]\mapsto X^{NI[m]}$, where $NI[m]$ is the nerve of the contractible groupoid with $(m+1)$ objects. Indeed, this follows by adjunction from the fact that $[m]\mapsto NI[m]$ defines a Reedy cofibrant cosimplicial diagram in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ that is homotopically constant. One can use this canonical simplicial resolution to explicitly compute the homotopy limit of a diagram of Segal spaces $X\colon\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, which does not need to be injectively fibrant. This can be expressed as a weighted limit, via the Bousfield–Kan formula [AØ23]: writing $X^{K}=\lim_{[m]\in(\Delta_{/K})^{\mathrm{op}}}X^{NI[m]}$, we have that

### v2

@@ 5. Properties of the model structure
In this section, we show that the categorical model structure $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ constructed in Theorem˜3.1 has desirable properties. Specifically, in Section˜5.1, we prove that it is cartesian closed, and in Section˜5.2, we establish that it is left proper.
@@ 5.1. Cartesian closedness
The model category $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ from Theorem˜3.1 is cartesian closed.
@@
By Proposition˜3.8, the $\mathcal{I}$-cofibrations $f\colon X\to Y$ and $f^{\prime}\colon X^{\prime}\to Y^{\prime}$ are monomorphisms and there are sets $R$ and $R^{\prime}$ and weak equivalences $X_{0}\amalg R\xrightarrow{\simeq}Y_{0}$ and $X^{\prime}_{0}\amalg R^{\prime}\xrightarrow{\simeq}Y^{\prime}_{0}$ in $\mathrm{sSet}$ whose restrictions to $X_{0}$ and $X^{\prime}_{0}$ are $f_{0}$ and $f^{\prime}_{0}$, respectively. We show that the pushout-product map $f\,\square\,f^{\prime}$ is an $\mathcal{I}$-cofibration by verifying the conditions from Proposition˜3.8.
Since the maps $f$ and $f^{\prime}$ are monomorphisms, so is their pushout-product $f\,\square\,f^{\prime}$. Consider the following maps in $\mathrm{sSet}$
$\left(X_{0}\times(X^{\prime}_{0}\amalg R^{\prime})\amalg_{X_{0}\times X^{\prime}_{0}}(X_{0}\amalg R)\times X^{\prime}_{0}\right)\amalg R\times R^{\prime}$$(X_{0}\amalg R)\times(X^{\prime}_{0}\amalg R^{\prime})$$\left(X_{0}\times X^{\prime}_{0}\amalg X_{0}\times R^{\prime}\amalg R\times X^{\prime}_{0}\right)\amalg R\times R^{\prime}$$Y_{0}\times Y^{\prime}_{0}$$\left(X_{0}\times Y^{\prime}_{0}\amalg_{X_{0}\times X^{\prime}_{0}}Y_{0}\times X^{\prime}_{0}\right)\amalg R\times R^{\prime}$
$\cong$
$\cong$
$\simeq\simeq$
The top map is a weak equivalence in $\mathrm{sSet}$ as a map between homotopy pushouts induced from a diagram of weak equivalences, and the bottom map is a weak equivalence in $\mathrm{sSet}$ since this is cartesian closed. Hence, by 22-out-of-33, the right-hand map is a weak equivalence in $\mathrm{sSet}$ whose restriction to $X_{0}\times Y^{\prime}_{0}\amalg_{X_{0}\times X^{\prime}_{0}}Y_{0}\times X^{\prime}_{0}$ is $(f\,\square\,f^{\prime})_{0}$. This shows that $f\,\square\,f^{\prime}$ is an $\mathcal{I}$-cofibration. ∎
We now further show that, if one of the two $\mathcal{I}$-cofibrations is a $\mathcal{J}$-cofibration, then their pushout-product is in fact a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. By [Joy08, Proposition E.2.14], this is sufficient to conclude the proof of Theorem˜5.1.
@@
We first deal with the case where $f^{\prime}$ is a $\mathcal{J}_{0}$-cofibration, and hence a trivial cofibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ by Remark˜3.10. By Proposition˜3.8, the $\mathcal{I}$-cofibration $f\colon X\to Y$ is a monomorphism, and so a cofibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$. Since $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ is cartesian closed, we get that the pushout-product map $f\,\square\,f^{\prime}$ is weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$. By 22-out-of-33, a $\mathcal{J}_{0}$-fibrant replacement of $f\,\square\,f^{\prime}$ is also a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ between Segal spaces, and so a Dwyer–Kan equivalence by Example˜2.12. This shows that $f\,\square\,f^{\prime}$ is a weak equivalence.
To treat the general case, it is now enough to treat the case $f^{\prime}$ being $F[0]\hookrightarrow NI[1]$, since $\mathcal{J}$-cofibrations are generated by maps in $\mathcal{J}_{0}$ and $F[0]\hookrightarrow NI[1]$. Showing that the $\mathcal{I}$-cofibration $X\times NI[1]\amalg_{X}Y\to Y\times NI[1]$ is a weak equivalence is equivalent to showing that, for every $\mathcal{J}$-fibration $W\to Z$ between $\mathcal{J}$-fibrant objects, the induced map
@@
is an $\mathcal{I}$-fibration. We prove this by showing that it is both an isofibration and a Dwyer–Kan equivalence. Since $W\to Z$ is an isofibration between Segal spaces, then Lemma˜2.18 shows that the desired map is an isofibration. Moreover, by Lemmas˜2.18 and 3.16, the maps $W^{NI[1]}\to W$ and $Z^{NI[1]}\to Z$ are both isofibrations and Dwyer–Kan equivalences; hence they are $\mathcal{I}$-fibrations. Then, in the composite
@@
the second map is an $\mathcal{I}$-fibration as the pullback of $Z^{NI[1]}\to Z$ and the composite is a Dwyer–Kan equivalence. Hence, the first map is a Dwyer–Kan equivalence as well, by 22-out-of-33. ∎
@@ 5.2. Left properness
The model category $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ from Theorem˜3.1 is left proper.
@@
Let $i\colon A\to B$ be an $\mathcal{I}$-cofibration and $f\colon A\to C$ be a weak equivalence $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. Consider the pushout $g\colon B\to D$ of $f$ along $i$; we want to prove that $g$ is also a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. First note that, without loss of generality, we can assume that $C$ is a Segal space, as $\mathcal{J}$-cofibrations are closed under pushouts and weak equivalences in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ satisfy 22-out-of-33.
@@
We need to check that $qk$ is a weak equivalence. Note that $k$ is a $\mathcal{J}$-cofibration, and hence a weak equivalence, as it is the pushout of the $\mathcal{J}$-cofibration $j$. To see that $q$ is a weak equivalence, note that the pushout $i^{\prime}$ of $i$ is an $\mathcal{I}$-cofibration, so in particular a monomorphism by Proposition˜3.8. Since $f$ and $j$ are both weak equivalences in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ and $C$ is a Segal space, the 22-out-of-33 property implies that the $\mathcal{J}$-fibration $p$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ between Segal spaces, and hence a Dwyer–Kan equivalence. By Theorem˜4.2 (iii), the map $p$ is then a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. As the model structure $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ is left proper, the pushout $q$ of $p$ along the monomorphism $i^{\prime}$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. By Lemma˜4.3, the map $q$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, so that the composite $qk$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ as well. ∎
@@ 6. Applications
In this last section, we present two notable consequences of the existence of the categorical model structure. In Section˜6.1, we show that the inclusion of categories into $(\infty,1)$-categories has a particularly simple description in terms of the model category $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. Specifically, we prove that the discrete nerve functor is right Quillen. Then, in Section˜6.2, we show that homotopy limits of $(\infty,1)$-categories can be computed in a manner similar to homotopy limits of ordinary categories. Specifically, we prove that the limit of an injectively fibrant diagram for $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ is already a homotopy limit for the complete Segal space model structure.
@@ 6.1. Nerves of categories
The adjunction from Remark˜2.7
$c\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\rightleftarrows\mathrm{Cat}\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0muN$
@@
Any map of sets is a fibration in $\mathrm{sSet}$, so that any map between simplicial sets (viewed as simplicial spaces) is a Reedy fibration. Since $N\mathcal{C}$ satisfies the Segal conditions, for every category $\mathcal{C}$, and the nerve preserves isofibrations, it follows that $N$ preserves fibrations. Next, the trivial fibrations $f\colon\mathcal{C}\to\mathcal{D}$ in $\mathrm{Cat}$ are the functors that are fully faithful and surjective on objects, so that the induced map $Nf\colon N\mathcal{C}\to N\mathcal{D}$ between the nerves satisfies the equivalent conditions of Lemma˜2.11. By Proposition˜3.17, the map $Nf$ is then an $\mathcal{I}$-fibration. This implies that $N$ is a right Quillen functor.
Finally, the nerve $N\mathcal{C}$ of any category $\mathcal{C}$ is cofibrant by Proposition˜3.8, so that the derived counit is equivalent to the counit. The latter is an isomorphism as the nerve is fully faithful. ∎
@@ 6.2. Homotopy limits of $(\infty,1)$-categories
In fact, using [Hir03, Proposition 19.9.4], this is a special instance of the following result for more general homotopy limits.
@@
Let $\mathcal{C}$ be a small category and consider a diagram $F\colon\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}$. Then any homotopy limit of $F$ in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ is a homotopy limit in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$.
@@
Consider an injective fibrant replacement $F\Rightarrow\widehat{F}$ in the model structure of diagrams $\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, and an injective fibrant replacement $\widehat{F}\Rightarrow\widetilde{F}$ in the model structure of diagrams $\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. For every object $c\in\mathcal{C}$, the induced map $\widehat{F}(c)\to\widetilde{F}(c)$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ between Segal spaces, and so a Dwyer–Kan equivalence by Theorem˜4.2 (iii). As $\mathrm{id}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ is a right Quillen functor, the diagram $\widetilde{F}\colon\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ is also injectively fibrant as a diagram $\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. Hence, we have a levelwise weak equivalence $\widehat{F}\Rightarrow\widetilde{F}$ between injectively fibrant diagrams in the model structure of diagrams $\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. As a consequence, the induced map $\mathrm{lim}_{\mathcal{C}}\widehat{F}\to\mathrm{lim}_{\mathcal{C}}\widetilde{F}$ is a Dwyer-Kan equivalence between Segal spaces, and so a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ by Theorem˜4.2 (iii). This concludes the proof. ∎
@@
Since $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ is cartesian closed by Theorem˜5.1, every Segal space $X$ has a canonical Reedy fibrant simplicial resolution $\Delta^{\mathrm{op}}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ sending $m]\mapsto X^{NI[m]}$, where $NI[m]$ is the nerve of the contractible groupoid with $(m+1)$ objects. Indeed, this follows by adjunction from the fact that $m]\mapsto NI[m]$ defines a Reedy cofibrant cosimplicial diagram in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ that is homotopically constant. One can use this canonical simplicial resolution to explicitly compute the homotopy limit of a diagram of Segal spaces $X\colon\mathcal{C}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, which does not need to be injectively fibrant. This can be expressed as a weighted limit, via the Bousfield–Kan formula [AØ23]: writing $X^{K}=\lim_{[m]\in(\Delta_{/K})^{\mathrm{op}}}X^{NI[m]}$, we have that